OpenStudy (anonymous):

Unit 6 Rational Expression Identify the restrictions on the domain. x - 7 / x x + 4 x + 3 A. x doesnt equal - 4, -3 B. x doesnt equal 7, 0 C. x doesnt equal -4, 0 D. x doesnt equal 7, -3

OpenStudy (triciaal):

basically you never want to multiply or divide by zero so the restriction is the value that would create that situation.

OpenStudy (triciaal):

hope your visits were fine. the original is not clearly written. can you use the "draw" below and write it out ?

OpenStudy (anonymous):

\[\frac{ x - 7 }{ x + 4} \div \frac{ x }{ x +3 }\]

OpenStudy (anonymous):

\[\frac{ x-7 }{ x+4 }\times \frac{ x+3 }{ x }\] restrictions are\[x+4\neq0,x \neq-4~and~x \neq0\]

OpenStudy (anonymous):

okay thanks I need it I have more question on the way